How to calculate the maximum area of rectangle give the fixed perimeter of 24.
Found 3 solutions by Boreal, ikleyn, Alan3354:
Answer by Boreal(15235)

length is x and width is 12-x, and that gives half perimeter of 12
area is x(12-x)
maximize their product, which is 12x-x^2
the vertex is at -b/2a or -12/-2 or x=6
12-x=6
This is a square with area 36 sq units
Calculus gets the same answer.

Answer by ikleyn(52787)

.

Maximum area of a rectangle with given perimeter is the area of a SQUARE with this perimeter.

It is the general rule.

See the lessons
    - HOW TO complete the square to find the minimum/maximum of a quadratic function
    - Briefly on finding the minimum/maximum of a quadratic function
    - HOW TO complete the square to find the vertex of a parabola
    - Briefly on finding the vertex of a parabola
    - A rectangle with a given perimeter which has the maximal area is a square
in this site.

So, in your case the perimeter is 24 units.

The square with such a perimeter has the side of 24/4 = 6 units,
and its area is 6*6 = 36 square units.

It is MAXIMAL possible area for rectangles with the perimeter of 24 units.


Answer by Alan3354(69443)

How to calculate the maximum area of rectangle give the fixed perimeter of 24.
------------
P = 2W + 2L = 24
W + L = 12
-----------------
1st try W = L = 6 and 6
Area = 6*6 = 36
Any changes --> Area = (W + c)*(W - c) = W^2 - c^2
Any value of c <> 0 ---> a smaller area.